Instruction for Making L^{A}TEX Compuscripts Using PTPTEX

Shin-Ichiro Tomonaga^{∗)} and Hideki Yukawa^{∗,}^{∗∗)} Physics Department, Tokyo Bunrika University, Tokyo 113

∗Yukawa Institute for Theoretical Physics, Kyoto University, Kyoto 606-8502

(Received )

Write your ABSTRACT here.

§1. Introduction

Start your paper from here.

§2. Equations

2.1. Fractions

§3. References Acknowledgements

We would like to thank ........... Appendix A First Appendix Appendix B Second Appendix

References

1) Y. Nambu and G. Jona-Lasinio, Phys. Rev. 122 (1961), 345.

M. Kobayashi and T. Maskawa, Prog. Theor. Phys. 49 (1973), 652.

2) A. Polyakov and P.B. Wiegmann, Phys. Lett. B131 (1983), 121.

P.B. Wiegmann, Phys. Lett. B141 (1984), 217; B142 (1984), 173.

3) R. Streater and A.S. Wightman, PCT, Spin & Statistics, and All That (W.A. Benjamin, New York, 1964), p. 117.

∗) A friend of Schwinger because of bearing ‘swing’ commonly in their names. ∗∗) A common friend of Fermi and Bose. E-mail address: yukawa@yukawa.kyoto-u.ac.jp

typeset using PTPTEX.sty <ver.1.0>